– Unlimited Distribution –

The hitec class

Eli Billauer The Company, Ltd

October 17, 2001

1 The general idea

This short paper is mainly a demonstration page to show what the papers written with this class will look like.

The document class is a hack on the well-known article class, where pieces of TEX code has been stolen from a couple of other classes.

This class completes, in my opinion, the set of tools needed to use L^{A}TEX in a hi-tec environment, where Microsoft rules too often. I’m delighted by the fact that Xemacs and L^{A}TEX run in a win32 environment, and that proper pdf’s can be produced using utilities such as pdflatex or dvipdfm. These allow me to get a Linux-feeling even when running Windows. Even better: I can write technical documents the way I like, and produce pdf’s that noone will complain about.

But I discovered that there was no way to escape the academic look of the wellknown article document class. There are many other classes around, yes, but almost all of them smell quite the same. Academy is not a four-letter word, but when a paper is submitted to your boss, he better not get the wrong impression before even reading it.

So, in order to solve my problem of giving my documents a nice outfit, I started the adventure of modifying article. While doing that, I understood why noone else has yet published such a class...

2 hitec vs. article

Papers that were compiled neatly as article are expected to give a fairly nice result right away after changing to hitec . The only exception is that hitec doesn’t work with two-column documents. Such documents’ compilation will be aborted with an error message.

Aside from that, the hitec class behaves a bit differently regarding document information:

Eli Billauer  – Unlimited Distribution –  The Company, Ltd

• The \author command can not be repeated to present more than one author. This wasn’t an attempt to reflect the loneliness of hi-tec workers, as opposed to the academic world, where they always work in pairs. The reason is technical: The author’s name appears on all pages (hi-tec, right?) and there’s no place for a list of people. In the case of multiple authors (did you waste time working together?) write them all in a single line. Remember that omitting an author can be a very painful mistake, and this will happen with no warning when switching to hitec .

• The \company command allows you not only to tell who you are, but also who you’re working for.

• The \confidential command has been added, to allow companies to mark their papers as confidential.

3 Just a small tip

If a pdf is your final target, going \usepackage{hyperref} in the beginning of your document is very recommended. The hyperref package will not only create weblike links where there are references to equations or sections, but it also creates the well-known bookmark list.

4 Summary

The hitec class was designed to allow L^{A}TEX to produce papers that suite the hi-tec world, in functionality and appearance. Together with other free software, a Windowsrunning PC can become a comfortable platform for creating impressive pdf documents using well-tested tools.